# VICON: A Foundation Model for Multi-Physics Fluid Dynamics via Vision In-Context Operator Networks

### v2

@@ Abstract
In-Context Operator Networks (ICONs) are models that learn operators across different types of PDEs using a few-shot, in-context approach. Although they show successful generalization to various PDEs, existing methods treat each data point as a single token, and suffer from computational inefficiency when processing dense data, limiting their application in higher spatial dimensions. In this work, we propose Vision In-Context Operator Networks (VICON), incorporating a vision transformer architecture that efficiently processes 2D functions through patch-wise operations. We evaluated our method on three fluid dynamics datasets, demonstrating both superior performance (reducing the rescaled $L^{2}$ error by $40\%$ and $61.6\%$ for two benchmark datasets for compressible flows, respectively) and computational efficiency (requiring only one-third of the inference time per frame) in long-term rollout predictions compared to the current state-of-the-art sequence-to-sequence model with fixed timestep prediction: Multiple Physics Pretraining (MPP). Compared to MPP, our method preserves the benefits of in-context operator learning, enabling flexible context formation when dealing with insufficient frame counts or varying timestep values.
@@ 1 Introduction
Machine learning has recently emerged as a powerful tool for solving Partial Differential Equations (PDEs). Traditional approaches primarily operate on discrete representations, with Convolutional Neural Networks (CNNs) excelling at processing regular grids (Gupta & Brandstetter, 2022) and Graph Neural Networks (GNNs) handling unstructured meshes (Sanchez-Gonzalez et al., 2020; Lino et al., 2022; Cao et al., 2023). Although recent developments have incorporated transformer architectures to capture global dependencies (Cao, 2021), these discrete approaches often face two key limitations: they lack explicit PDE context information and are confined to operating on fixed or similar topologies.
Several methods have emerged to address these limitations. The Deep Galerkin Method (DGM) (Sirignano & Spiliopoulos, 2018), Deep Ritz Method (DRM) (E & Yu, 2018), and Physics-Informed Neural Networks (PINN) (Raissi et al., 2019) incorporate PDE constraints in the model loss terms. The model, once trained, becomes a surrogate solution that can be queried at any location in the domain. However, these methods typically require complete retraining when initial or boundary conditions change. This limitation led to the development of “operator learning,” where neural networks map input functions to output functions, such as from initial conditions to later-time solutions. This capability was first demonstrated using shallow neural networks (Chen & Chen, 1995b, a), followed by specialized architectures such as the widely-adopted Fourier Neural Operator (FNO) (Li et al., 2021; Kovachki et al., 2023), Deep Operator Network (DeepONet) (Lu et al., 2021), and recent attempts to incorporate transformers as operator learners (Li et al., 2022).
Although the aforementioned methods excel at learning individual, single operators, they are not able to generalize across different operators. This limitation necessitates costly retraining for new equation parameters or predicting with longer time horizons. Inspired by the success of large language models (LLMs) in multi-domain generation, recent work has explored “PDE foundation models”. Common approaches employ pre-training and fine-tuning strategies (Chen et al., 2024; Herde et al., 2024; Zhang et al., 2024c), though these require additional deployment resources. To achieve zero-shot PDE foundation models, additional information is often needed, for example, the PROSE approach (Liu et al., 2024b, a; Sun et al., 2024a; Jollie et al., 2024) encodes the governing equations directly through symbolic information. Other approaches achieve zero- or few-shot generalization by using directed graphs to represent equations (Ye et al., 2024) or symbol tokenization (Lorsung et al., 2024).
@@
While these approaches require explicit encoding of PDE information, which may not be readily available or precisely known in practice, the In-Context Operator Network (ICON) (Yang et al., 2023a) offers a fundamentally different approach. ICON implicitly encodes system dynamics through input-output function pairs (before-after operators), achieving a few-shot generalization without retraining or explicit PDE knowledge. A single ICON model has shown success in handling both forward and inverse problems in various ODEs, PDEs, and mean field control scenarios.
@@
However, ICON’s computational efficiency becomes a critical bottleneck when handling dense data in higher dimensions. Specifically, representing each spatial point as an individual token leads to quadratic complexity in the number of points, making it computationally prohibitive for practical 2D and 3D applications - to date, only one instance of a 2D case (Yang et al., 2023a) using sparse data points has been demonstrated.
@@
To address this limitation, we propose Vision In-Context Operator Networks (VICON), which leverage vision transformers to process 2D functions using an efficient patch-wise approach. Going beyond classic vision transformers, which typically process patches from a single image, VICON extends the architecture to handle sequences of input-output function pairs, allowing “next function prediction” capabilities while preserving the benefits of in-context operator learning. Our contributions include:
@@
  • Development of VICON, a vision transformer-based in-context operator network for two-dimensional time-dependent PDEs that maintains the in-context operator learning framework and handles dense data points in higher dimensions.
  • We evaluated our model across diverse fluid dynamics systems, comprising 879K frames with varying timestep sizes. Empirical results demonstrate significant improvements over the state-of-the-art MPP (McCabe et al., 2023), including a 40% and 61. 6% reduction in the scaled L2 error for long-term predictions in two benchmark datasets, a 67% reduction in the prediction error of turbulence kinetic energy, and approximately one third of the singe step inference time.
  • VICON enables flexible rollout scheme unattainable by traditional sequence-to-sequence models: it can form example pairs from nonsequential time sequences, handle scenarios with insufficient example frames, and adapt to varying timestep strides without retraining. These capabilities allow for robust rollout even with incomplete or irregularly sampled temporal data.
@@ 2 Related Works
Operator learning (Chen & Chen, 1995b, a; Li et al., 2021; Lu et al., 2021) addresses the challenge of approximating operators $G:U\to V$, where $U,V$ are function spaces. These operators usually model solutions of physical systems and differential equations, representing the map from initial/boundary conditions and parameters to equation solutions. DeepONets (Lu et al., 2021, 2022) use branch and trunk networks to separately process inputs and query points, and have been extended to handle multiple input parametric functions in MIONet (Jin et al., 2022). FNOs (Li et al., 2021) leverage kernel integration approximations for PDE solutions by processing features in Fourier space, where kernel integrations can be efficiently computed through multiplications. These methods have been further extended to incorporate equation information (Wang et al., 2021), multiscale features (Zhang et al., 2024a; Wen et al., 2022), different input/output meshes (Zhang et al., 2023), complex geometries (Li et al., 2024), and distributed learning (Zhang et al., 2024b). The applications of operator learning include weather prediction (Pathak et al., 2022; Jiang et al., 2024), biology (Yin et al., 2024), and uncertainty quantification (Moya et al., 2024). However, these approaches are typically designed to learn a single operator, requiring retraining when encountering different types of PDE or time-step sizes.
PDE Foundation Models
Foundation models are large-scale, pre-trained models that serve as versatile building blocks for a wide range of downstream tasks in natural language processing (Brown et al., 2020; Touvron et al., 2023), computer vision (Ramesh et al., 2021), and other AI applications. However, they are not directly applicable to scientific computing tasks such as the discovery of PDEs (Schaeffer, 2017) or computational fluid dynamics (Wang et al., 2024), where high precision is essential. Recent work has explored adapting the principles of the foundation model to PDEs through various approaches. Some methods employ pre-training and fine-tuning strategies (Chen et al., 2024; Herde et al., 2024), though this requires additional computational resources during deployment. Zero-shot PDE foundation models incorporate additional information; for example, the PROSE approach (Liu et al., 2024b, a; Sun et al., 2024a; Jollie et al., 2024) encodes the PDE directly through symbolic information, which can also be fine-tuned (Sun et al., 2024b). Other models achieve zero- or few-shot generalization by adding PDE structures to the network (Lorsung et al., 2024; Ye et al., 2024); however, these methods require prior knowledge of the PDE structure and information on the data.
In-Context Operator Networks
@@
ICONs (Yang et al., 2023a, b; Yang & Osher, 2024) learn operators through observing input-output function pairs, enabling few-shot generalization across various PDEs without explicit PDE representation or fine-tuning. ICONs have demonstrated success in handling forward and inverse problems in ODEs, PDEs, and mean-field control scenarios. However, they face significant computational challenges when processing dense data, as they represent functions through scattered point tokens, resulting in quadratic computational complexity. This limitation has largely restricted their application to 1D problems, with only sparse sampling possible in 2D cases. Our work addresses these limitations by introducing a vision transformer architecture that efficiently handles dense 2D data while maintaining the benefits of in-context operator learning.
@@
Figure 1: VICON model overview. The inputs to the model are pairs of conditions (COND) and quantities of interest (QOI), which are patchified and flattened before feeding into the transformer layers. The outputs, which represent different patches in the output frame, are transformed back to obtain the final predictions.
@@ 3 Preliminaries
ICONs were introduced and developed in (Yang et al., 2023a, b; Yang & Osher, 2024). They adapt the in-context learning framework from large language models, aiming to train a model that can learn operators from prompted function examples.
@@
For training parallelization, ICON uses a special causal attention mask to perform autoregressive learning (i.e., the model only sees the leading $J$ pairs and ${\bm{c}}_{J+1}$ when predicting $\tilde{{\bm{q}}}_{J+1}$) and enables output of all predictions within a single forward pass (Yang et al., 2023b):
@@
Training loss is computed as the mean squared error (MSE) between the predicted states $\tilde{{\bm{q}}}_{i}$ and the ground truth states ${\bm{q}}_{i}$. To ensure that the model receives sufficient contextual information about the underlying dynamics, we only compute errors for the indices $i>I_{\text{min}}$, effectively requiring at least $I_{\text{min}}$ examples in context. This approach has demonstrated empirical improvements in model performance.
@@
where $J\in\{I_{\text{min}},\dots,I-1\}$ is the number of in-context examples, and $k>I$ is the index of the condition for future prediction.
@@
Table 1: (Comparison) Summary of Rollout Rescaled $L^{2}$ Error (scale by std) for different methods across various cases. The best results are highlighted in bold.
Rollout Rescaled $L^{2}$ Error Case Ours (single step) Ours (multi step) MPP (McCabe et al., 2023)
Step 1 [1e-2] PDEArena-Incomp 11.10 11.10 6.17
PDEBench-Comp-LowVis 15.61 15.61 16.37
PDEBench-Comp-HighVis 5.97 5.97 20.90
Step 5 [1e-2] PDEArena-Incomp 23.00 25.79 14.81
PDEBench-Comp-LowVis 24.56 39.54 45.45
PDEBench-Comp-HighVis 19.73 28.24 42.93
Step 10 [1e-2] PDEArena-Incomp 36.18 35.02 27.89
PDEBench-Comp-LowVis 37.47 51.50 64.11
PDEBench-Comp-HighVis 57.88 79.52 144.46
Last Step [1e-2] PDEArena-Incomp 77.81 70.83 93.52
PDEBench-Comp-LowVis 39.03 48.82 65.32
PDEBench-Comp-HighVis 71.17 96.23 185.29
All average [1e-2] PDEArena-Incomp 56.27 51.32 55.95
PDEBench-Comp-LowVis 27.08 36.29 46.68
PDEBench-Comp-HighVis 30.06 40.06 72.37
@@ 4.1 Problem Setup
Table 2: Summary of Rollout Rescaled $L^{2}$ Error Metrics (single step, scale by std) for Ablation Studies across various cases. The best results are highlighted in bold. Columns: N = PDEArena-Incomp only, E = PDEBench-Comp-LowVis only, C = PDEBench-Comp-HighVis only, N + E = PDEArena-Incomp + PDEBench-Comp-LowVis, N + C = PDEArena-Incomp + PDEBench-Comp-HighVis, E + C = PDEBench-Comp-LowVis + PDEBench-Comp-HighVis. All ablation groups have equal dataset batch sizes, adding up to 30. The best results are marked with bold, 2nd-best with underline.
Rollout Rescaled $L^{2}$ Error [1e-2] Case All included N E C N + E N + C E + C
Step 1 PDEArena-Incomp 11.10 6.72 54.26 58.63 11.11 7.84 53.69
PDEBench-Comp-LowVis 15.61 88.24 13.59 46.66 14.86 45.96 14.08
PDEBench-Comp-HighVis 5.79 216.50 49.68 2.81 38.38 3.28 4.65
Step 5 PDEArena-Incomp 23.00 14.18 105.19 85.76 22.95 16.25 100.12
PDEBench-Comp-LowVis 24.56 106.29 22.10 67.27 23.88 68.15 22.53
PDEBench-Comp-HighVis 19.73 584.50 170.20 7.49 141.22 8.38 12.41
Step 10 PDEArena-Incomp 36.18 23.90 147.28 110.04 36.87 26.63 142.60
PDEBench-Comp-LowVis 37.47 119.84 33.94 81.27 36.43 82.27 34.49
PDEBench-Comp-HighVis 57.88 1903.41 514.88 20.26 410.23 22.65 33.42
Last Step PDEArena-Incomp 77.81 69.95 289.62 185.34 86.16 71.90 294.19
PDEBench-Comp-LowVis 39.03 120.74 35.16 83.24 37.61 84.19 35.70
PDEBench-Comp-HighVis 71.17 2356.29 626.60 24.35 495.02 27.28 40.24
All average PDEArena-Incomp 56.26 44.35 202.67 142.00 59.65 47.31 203.04
PDEBench-Comp-LowVis 27.08 107.54 24.35 68.71 26.27 69.52 24.83
PDEBench-Comp-HighVis 30.06 956.67 263.98 10.91 213.34 12.25 18.06
@@ 4.2 Vision In-Context Operator Networks
Inspired by the Vision Transformer (ViT) (Dosovitskiy, 2020), we address these limitations by dividing the physical fields into patches, where each patch is flattened and projected as a token. This approach, which we call Vision In-Context Operator Networks (VICON), has its forward process illustrated in Figure 1 and detailed in the following.
@@
$\hat{{\bm{c}}}_{i}^{k}$ $=f_{\phi}({\bm{C}}_{i}^{k}),$ (4)
$\hat{{\bm{q}}}_{i}^{l}$ $=f_{\phi}({\bm{Q}}_{i}^{l}),$
where $k=1,\ldots,N_{c}$ and $l=1,\ldots,N_{q}$ index the patches.
@@
${\bm{c}}_{i}^{k}$ $=\hat{{\bm{c}}}_{i}^{k}+\mathbf{E}_{p}(k)+\mathbf{E}_{c}(i),$ (5)
${\bm{q}}_{i}^{l}$ $=\hat{{\bm{q}}}_{i}^{l}+\mathbf{E}_{p}(l)+\mathbf{E}_{q}(i).$
@@
To support autoregressive prediction similar to Eq 2, VICON employs an alternating-sized (in the case where $N_{c}\neq N_{q}$) block causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models (Brown et al., 2020) and large vision models (Bai et al., 2024). The mask is defined as follows:
$\mathbf{M}=\small\begin{bmatrix}\mathbf{1}_{N_{c},N_{c}}&\mathbf{0}&\cdots&\mathbf{0}&\mathbf{0}\\ \mathbf{1}_{N_{q},N_{c}}&\mathbf{1}_{N_{q},N_{q}}&\cdots&\mathbf{0}&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \mathbf{1}_{N_{c},N_{c}}&\mathbf{1}_{N_{c},N_{q}}&\cdots&\mathbf{1}_{N_{c},N_{c}}&\mathbf{0}\\ \mathbf{1}_{N_{q},N_{c}}&\mathbf{1}_{N_{q},N_{q}}&\cdots&\mathbf{1}_{N_{q},N_{c}}&\mathbf{1}_{N_{q},N_{q}}\end{bmatrix}$ (6)
@@
After obtaining the output tokens from Eq 2, we extract the tokens corresponding to the input patch indices $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@
Throughout our experiments, we set $I_{\text{min}}=5$ in-context example pairs, providing at least five examples to ensure sufficient information about the dynamics before evaluating errors.

### v3

# VICON: Vision In-Context Operator Networks for Multi-Physics Fluid Dynamics Prediction

@@ Abstract
In-Context Operator Networks (ICONs) have demonstrated the ability to learn operators across diverse partial differential equations using few-shot, in-context learning. However, existing ICONs process each spatial point as an individual token, severely limiting computational efficiency when handling dense data in higher spatial dimensions. We propose Vision In-Context Operator Networks (VICON), which integrates vision transformer architectures to efficiently process 2D data through patch-wise operations while preserving ICON’s adaptability to multiphysics systems and varying timesteps. Evaluated across three fluid dynamics benchmarks, VICON significantly outperforms state-of-the-art baselines: DPOT and MPP, reducing the averaged last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times. VICON naturally supports flexible rollout strategies with varying timestep strides, enabling immediate deployment in imperfect measurement systems where sampling frequencies may differ or frames might be dropped—common challenges in real-world settings—without requiring retraining or interpolation. In these realistic scenarios, VICON exhibits remarkable robustness, experiencing only 24.41% relative performance degradation compared to 71.37%-74.49% degradation in baseline methods, demonstrating its versatility for depolying in realistic applications. Our scripts for processing datasets and code are publicly available at https://github.com/Eydcao/VICON.
@@ 1 Introduction
Machine learning has emerged as a powerful tool for solving Partial Differential Equations (PDEs). Traditional approaches primarily operate on discrete representations, with Convolutional Neural Networks (CNNs) excelling at processing regular grids [1], Graph Neural Networks (GNNs) handling unstructured meshes [2, 3, 4], and transformers capturing dependencies between sampling points within the same domains [5, 6, 7]. However, these discrete approaches face a fundamental limitation: they lack PDE context information (such as governing parameters) and consequently cannot generalize to new parameters or PDEs beyond their training set.
This limitation led to the development of “operator learning,” where neural networks learn mappings from input functions to output functions, such as from initial/boundary conditions to PDE solutions. This capability was first demonstrated using shallow neural networks [8, 9], followed by specialized architectures including the widely-adopted Deep Operator Network (DeepONet) [10], Fourier Neural Operator (FNO) [11, 12], and recent transformer-based operator learners [13, 14, 15]. While these methods excel at learning individual parametric PDEs, they cannot generalize across different PDE types, necessitating costly retraining for new equations or even different time step sizes.
Inspired by the success of large language models in multi-domain generation, researchers have explored “Multi-Physics PDE models” to address this single-operator limitation. Common approaches employ pre-training and fine-tuning strategies [16, 17, 18], though these require substantial additional data (typically hundreds of frames) and fine-tuning before deployment. This requirement severely limits their practical applications in scenarios like online control and data assimilation, where rapid adaptation with limited data sampling is essential.
To eliminate the need for additional data collection and fine-tuning, several works have attempted to achieve few-shot or zero-shot generalization across multi-physics systems by incorporating explicit PDE information. For example, the PROSE approach [19, 20, 21, 22] processes the symbolic form of governing equations using an additional transformer branch. Similarly, PDEformer [23] converts PDEs into directed graphs as input for graph transformers, while others embed symbol-tokens directly into the model [24]. Despite these advances, such methods require explicit knowledge of the underlying PDEs, which is often unavailable when deploying models in new environments.
@@
To further lower barriers for applying multi-physics models in real-world applications, the In-Context Operator Network (ICON) [25] offers a fundamentally different approach: ICON implicitly encodes system dynamics through a few input-output function pairs, then extracts dynamics from these pairs in an in-context fashion. This approach enables few-shot generalization without retraining or explicit PDE knowledge, while the minimal required input-output pairs can be readily collected during deployment. A single ICON model has demonstrated success in handling both forward and inverse problems across various ODEs, PDEs, and mean field control scenarios.
@@
However, ICON’s computational efficiency becomes a critical bottleneck when handling dense data in higher dimensions. Specifically, ICON treats each sampled spatial point as an individual token, leading to quadratic computational complexity with respect to the number of points. This makes it computationally prohibitive for practical 2D and 3D applications—to date, only one instance of a 2D case using sparse data points has been demonstrated [25].
@@
To address this limitation, we propose Vision In-Context Operator Networks (VICON), which leverage vision transformers to process 2D functions using an efficient patch-wise approach. Going beyond classic vision transformers that typically process patches from a single image, VICON extends the architecture to handle sequences of input-output function pairs, enabling “next function prediction” capabilities while preserving the benefits of in-context operator learning.
Our contributions include:
@@
  • Development of VICON, a vision transformer-based in-context operator network for two-dimensional time-dependent PDEs that maintains the flexibility of in-context operator learning while efficiently handling dense data in higher dimensions.
  • Comprehensive evaluation across diverse fluid dynamics systems comprising 879K frames with varying timestep sizes, demonstrating substantial improvements over state-of-the-art models MPP [26] and DPOT [27] in both accuracy and efficiency. Our approach reduces the averaged last-step rollout error by 37.9% compared to DPOT and 44.7% compared to MPP, while requiring only 72.5% and 34.8% of their respective inference times.
  • Enhanced flexibility supporting varying timestep strides and non-sequential measurements, offering substantial robustness in realistic scenarios with imperfect data collection, e.g. with missing frames. Under these scenarios, VICON exhibits only 24.41% relative performance degradation versus 71.37%-74.49% degradation in baseline methods.
@@ 2 Related Works
Operator learning [8, 9, 11, 10] addresses the challenge of approximating operators $G:U\to V$, where $U$ and $V$ are function spaces representing physical systems and differential equations, e.g., $G$ maps initial/boundary conditions or system parameters to corresponding solutions. Among the various operator learning approaches, DeepONets [10, 28] employ branch and trunk networks to independently process inputs and query points, while FNOs [11] leverage fast Fourier transforms to efficiently compute kernel integrations for PDE solutions in regular domains. These methods have been extended to incorporate equation information [29, 30], multiscale features [31, 32, 33], adaptation to heterogeneous and irregular meshes [34, 35, 36, 37].
However, these approaches typically learn a single operator, requiring retraining when encountering different PDE types or time-step sizes. Our work aims to develop a unified model for multiple PDEs with few-shot generalization capabilities based on limited observation frames.
Multi-Physics PDE Models.
Foundation models in natural language processing [38, 39] and computer vision [40] have demonstrated remarkable versatility across diverse tasks. Drawing inspiration from this paradigm, recent research has explored developing unified models for multiple PDEs in domains such as PDE discovery [41] and computational fluid dynamics [42]. Several approaches follow pre-training and fine-tuning strategies [16, 17], though these require additional resources for data acquisition and fine-tuning before deployment. In parallel, researchers have developed zero- or few-shot multi-physics models by incorporating additional information; for instance, the PROSE approach [19, 20, 21, 22, 43] directly encodes symbolic information of PDEs into the model. Other methods achieving zero- or few-shot generalization include using physics-informed tokens [24], representing PDE structures as graphs [23], or conditioning transformers on PDE descriptions [44]; however, all these techniques require prior knowledge of the PDEs as additional inputs. We select two state-of-the-art models, MPP [26] and DPOT [27], as our baselines.
In-Context Operator Networks.
@@
ICONs [25, 45, 46] learn operators by observing input-output function pairs, enabling few-shot generalization across various PDEs without requiring explicit PDE representations or fine-tuning. These networks have demonstrated success in handling forward and inverse problems for ODEs, PDEs, and mean-field control scenarios. However, ICONs face significant computational challenges when processing dense data, as they represent functions through scattered point tokens, resulting in quadratic computational complexity with respect to the number of sample points. This limitation has largely restricted their application to 1D problems, with only sparse sampling feasible in 2D cases [25]. Our work addresses these limitations by introducing a vision transformer architecture that efficiently handles dense 2D data while preserving the benefits of in-context operator learning.
@@
Figure 1: VICON model overview. (a) The formation process for conditions (COND) and quantities of interest (QOI) pairs. $\Delta t$ is randomly sampled during training. (b) Model illustration. The inputs to the model are pairs of COND and QOI, which are patchified and flattened before feeding into the transformer layers. The outputs, which represent different patches in the output frame, are transformed back to obtain the final predictions. (c) With imperfect temporal measurements, VICON forms pairs using only clean data, and does not need to fill missing frames.
@@ 3 Preliminaries
ICONs were introduced and developed in [25, 45, 46]. They adapt the in-context learning framework from large language models, aiming to train a model that can learn operators from prompted function examples.
@@
For training parallelization, ICON uses a special causal attention mask to perform autoregressive learning (i.e., the model only sees the leading $J$ pairs and ${\bm{c}}_{J+1}$ when predicting $\tilde{{\bm{q}}}_{J+1}$) and enables output of all predictions within a single forward pass [45]:
@@
Training loss is computed as the mean squared error (MSE) between the predicted states $\tilde{{\bm{q}}}_{i}$ and the ground truth states ${\bm{q}}_{i}$. To ensure that the model receives sufficient contextual information about the underlying dynamics, we only compute errors for the indices $i>I_{\text{min}}$, effectively requiring at least $I_{\text{min}}$ examples in context. This approach has shown empirical improvements in model performance.
@@
where $J\in\{I_{\text{min}},\dots,I-1\}$ is the number of in-context examples, $I_{\text{min}}$ is the number of context exempted from loss calculation, and $k>I$ is the index of the condition for future prediction.
@@ 4.2 Vision In-Context Operator Networks
Inspired by the Vision Transformer (ViT) [47], we address these limitations by dividing the physical fields into patches, where each patch is flattened and projected as a token. This approach, which we call Vision In-Context Operator Networks (VICON), has its forward process illustrated in Figure 1(b) and detailed in the following.
@@
$\hat{{\bm{c}}}_{i}^{k}=f_{\phi}({\bm{C}}_{i}^{k}),\quad\hat{{\bm{q}}}_{i}^{l}=f_{\phi}({\bm{Q}}_{i}^{l}),$ (4)
where $k=1,\ldots,N_{c}\text{ and }l=1,\ldots,N_{q}\text{ are the index of patches.}$.
@@
${\bm{c}}_{i}^{k}=\hat{{\bm{c}}}_{i}^{k}+\mathbf{E}_{p}(k)+\mathbf{E}_{c}(i),\quad{\bm{q}}_{i}^{l}=\hat{{\bm{q}}}_{i}^{l}+\mathbf{E}_{p}(l)+\mathbf{E}_{q}(i).$ (5)
@@
To support autoregressive prediction similar to Equation 2, VICON employs an alternating-sized (in the case where $N_{c}\neq N_{q}$) block causal attention mask, as opposed to the conventional triangular causal attention mask as in mainstream generative large language models [38] and large vision models [48]. The mask is defined as follows:
$\mathbf{M}=\begin{bmatrix}\mathbf{1}_{N_{c},N_{c}}&\mathbf{0}&\cdots&\mathbf{0}&\mathbf{0}\\ \mathbf{1}_{N_{q},N_{c}}&\mathbf{1}_{N_{q},N_{q}}&\cdots&\mathbf{0}&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \mathbf{1}_{N_{c},N_{c}}&\mathbf{1}_{N_{c},N_{q}}&\cdots&\mathbf{1}_{N_{c},N_{c}}&\mathbf{0}\\ \mathbf{1}_{N_{q},N_{c}}&\mathbf{1}_{N_{q},N_{q}}&\cdots&\mathbf{1}_{N_{q},N_{c}}&\mathbf{1}_{N_{q},N_{q}}\end{bmatrix}$ (6)
@@
After obtaining the output tokens from Equation 2, we extract the tokens corresponding to the input patch indices $\mathbf{c}_{i}^{1},\ldots,\mathbf{c}_{i}^{N_{c}}$ and denote them as $\tilde{{\bm{q}}}_{i}^{1},\ldots,\tilde{{\bm{q}}}_{i}^{N_{c}}$. These tokens are then projected back to the original physical space using a shared learnable linear function $g_{\psi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{R_{x}\times R_{y}\times c}$:
@@
4.3 Prompt Normalization
$\mathcal{L}^{\prime}(\mathcal{N}_{c}(\mathbf{u}))=\mathcal{N}_{q}(\mathcal{L}(\mathbf{u})).$ (8)
4.4 Datasets and Data Augmentation
We evaluate on three fluid dynamics datasets representing different physical regimes: 1) PDEArena-Incomp [1] (incompressible Navier-Stokes equations), containing 2,496/608/608 trajectories (train/valid/test) with 56 timesteps each; 2) PDEBench-Comp-HighVis [49] (compressible Navier-Stokes with high viscosity), containing 40,000 trajectories of 21 timesteps each; and 3) PDEBench-Comp-LowVis [49] (compressible Navier-Stokes with numerically zero viscosity), containing 4,000 trajectories of 21 timesteps each. For PDEBench-Comp-HighVis and PDEBench-Comp-LowVis, we randomly split trajectories in 80%/10%/10% proportions for training/validation/testing, respectively.
The temporally homogeneous Markovian property (Section 4.1) enables natural data augmentation for training and flexible rollout strategy for inference (Section 4.5): by striding with larger timesteps to reduce the number of autoregressive steps in long-term prediction tasks: $\Delta t=s\Delta\tau_{i_{p}}\;|\;s=1,2,\dots,s_{\text{max}}$, where $\tau_{i_{p}}$ is the timestep size for recording in trajectory $i_{p}$. During training, for each trajectory in the training set, we first sample a stride size $s\sim\mathcal{U}\{1,\ldots,s_{\text{max}}\}$, then form $\langle\mathbf{u}_{t},\mathbf{u}_{t+s\Delta\tau}\rangle$ pairs for the corresponding operator. We illustrate this augmentation process in Figure 1(a), where different strides are randomly sampled during dataloading.
Additional dataset details appear in Appendix A.
4.5 Inference with Flexible Strategies
For long-term rollout, we employ VICON in an autoregressive fashion. The ability to predict with different timestep strides enables various rollout strategies. We explore two natural approaches: single-step rollout and flexible-step rollout, where the latter advances in larger strides to reduce the number of rollout steps.
For single-step rollout, we simply follow the autoregressive procedure with $s=1$. Flexible-step rollout involves a more sophisticated approach: given a maximum prediction stride $s_{\text{max}}$, we first make sequential predictions with a gradually growing $s=1$ to $s_{\text{max}}$ using $\mathbf{u}_{I_{0}}$ as the question condition, obtaining $\left\{\mathbf{u}_{I_{0}+s}\right\}_{s=1}^{s_{\text{max}}}$. Using each frame in this sequence as a question condition, we then make consistent $s_{\text{max}}$-step predictions while preserving all intermediate frames. This flexible-step strategy follows the approach in [46].
This strategy is also applicable to the imperfect measuring cases where partial input frames are missing, eg, due to sensor device error. In this case, we can still form pairs from the remaining frames, with minor adjustment to the strategy generation algorithms, as shown in Figure 1(c).
More details (including algorithms and examples) on rollout strategies are provided in Section C.1 (for perfect measurements) and in Section C.2 (for imperfect measurements).
4.6 Evaluation Metric
We evaluate the rollout accuracy of VICON using two different strategies described in Section 4.5. The evaluation uses relative and absolute $L^{2}$ errors between the predicted and ground truth frames, starting from the frame $I_{0}+1$. For relative scaling coefficients, we use channel-wise scaling standard deviation $\mathbf{\sigma}$ of ground truth frames, which vary between different prompts.
5 Experimental Results
We benchmark VICON against two state-of-the-art sequence-to-sequence models: DPOT [27] (122M parameters) and MPP [26] (AViT-B architecture, 116M parameters). The vanilla ICON, which would require processing over 114K tokens per frame (128×128×7), exceeds our available GPU memory and is thus computationally infeasible for direct comparison on these datasets.
For evaluation, we use both absolute and relative $L^{2}$ RMSE (Section 4.6) on rollout predictions. Since VICON offers the flexibility of predictions with different time step strides, we evaluate our single trained model using both single-step and flexible-step rollout strategies (Section 4.5) during inference.
For conciseness, we present summarized plots and tables the main text, and defer complete results and visualizations to Appendix D. Ablation studies examining key design choices of VICON—including patch resolution, positional encoding, and context length—are presented in Section D.1.
Figure 2: Main experiment results. (a) Last step rollout errors on 3 datasets. VICON outperforms MPP on all datasets and outperforms DPOT on 2 datasets. (b) VICON allows flexible rollout strategies to reduce error accumulation and demonstrates stride extrapolation. (c) VICON is robust to imperfect temporal measurements, while MPP and DPOT suffer from performance degradation. (d) VICON is smaller in size and has faster rollout time per step.
Table 1: Summary of Rollout Relative $L^{2}$ Error (scaled by std) across different methods and datasets. The best results are highlighted in bold. For flexible step rollout, step 3 works best for the PDEArena-Incomp dataset.
Rollout Relative $L^{2}$ Error Case Ours (single step) Ours (flexible step) DPOT MPP
Last step [1e-2] PDEArena-Incomp 76.77 68.03 65.27 93.52
PDEBench-Comp-LowVis 39.11 39.11 48.92 65.32
PDEBench-Comp-HighVis 61.41 61.41 2866 185.3
All average [1e-2] PDEArena-Incomp 56.27 48.50 41.20 55.95
PDEBench-Comp-LowVis 27.08 27.08 37.72 46.68
PDEBench-Comp-HighVis 30.06 30.06 821.9 72.37